How do you find the standard form of the equation of a circle, with a center (3,-1) through the point (-5,1)?
Answer by Alan3354(8317)

How do you find the standard form of the equation of a circle, with a center (3,-1) through the point (-5,1)?
--------------
Find the distance to the point, that's the radius.
r^2 = diffy^2 + diffx^2
r^2 = 2^2 + 8^2
r^2 = 68
------------
Then it's (x-h)^2 + (y-k)^2 = r^2 (h,k) is the center
(x-3)^2 + (y+1)^2 = 68
----------------
It can be expanded and rearranged.